# SkyMask: Attack-agnostic Robust Federated Learning with Fine-grained Learnable Masks

### v1

@@ III-C3 Mask clustering and classification
Figure 3: Visualizing the training process of masks applied to malicious and benign models, respectively.
III-D Why SkyMask Works
@@
We conduct an example experiment to show how SkyMask works. We choose a four-layer CNN model and Fashion-MNIST dataset and set 20% of 100 clients as malicious. Under Fang-Trim attack, we take the mask training process in one server-client communication round. We calculate the proportion of “1”s in each part of the binary mask, and a smaller value means fewer parameters of this client are selected for the aggregated masked model. For simplicity, here only shows 20 masks for malicious and 20 for benign clients in each mask training iteration.
@@
As Figure 3 shown, during the mask training process, the binary parameters in the “conv1.weight”, “conv2.weight” and “fc1.weight” parts of malicious model updates are set to zero more than those of benign updates. It is related to the fact that the attack prefers to attack the parameters in specific parts. The masks gradually reduce the poisoning effect of these toxic parameters during the training so that we can classify malicious clients with the help of these masks. Therefore, SkyMask can perceive the poisoning attacks at the parameter-level in the model updates without knowing the details of the attack, which makes it attack-agnostic.
III-E Complexity
@@
Let $O(T)$ denote the computational complexity of one local training iteration, and the model parameters’ dimensionality is $V$. To update the masks once, the server calculates the gradient of each mask from the gradient of the aggregated masked model $\tilde{W}_{t+1}$ by backpropagation, and its computational complexity is $O(nV)$.
@@
Hence, the total computational complexity is $O((T+nV)t_{m})+O(nV)=O(Tt_{m}+nVt_{m})$, where $t_{m}$ is the number of mask training convergence iterations. Since there is no data dependency between any mask parameters, we can use parallel training [21] to accelerate the mask optimization process so that the time overhead can be further reduced, trading computation power for time.
FLTrust only needs to calculate the similarity of local models and its baseline model, whose computational complexity is $O(nV)$. Trim sorts each parameter of local models and choose the middle part of the values, so its computational complexity is $O(nV\log n)$. Krum calculates the distance score of each local model, which is the sum of the distance between one local model and its $n-n_{m}-2$ nearest local models, so its computational complexity is $O(n^{2}V)$.
The above analysis shows that the additional cost of the mask learning mechanism over other defense methods is equivalent to some iterations of local training, which are acceptable compared to the computational power of the server.
@@
The space complexity of existing FL defense algorithms is either $O(n+nV)$ or $O(n^{2}+nV)$. Since we assign a mask for each client, the additional space complexity is also proportional to the number of clients $n$. It is proportional to the space complexity of the global model parameters because the masks have the same dimensionality as the global model parameters. Considering that the server stores $n$ local models’ parameters in the current round, the total space cost is about twice the others, and its space complexity is $O(nV+nV)=O(nV)$, which is reasonable.
IV Evaluation
@@
We implement a prototype of SkyMask based on PyTorch and evaluate our algorithm with existing Byzantine-robust FL aggregation algorithms and two malicious client detection algorithms. To verify the generalization of our method, we conduct experiments using several popular datasets and pair them with different models, respectively. We evaluate our approach with six popular attacks, explore the impact of the fraction of malicious clients, and verify its scalability. In sensitivity analysis, we discuss the influence of binarization threshold value, the number of features after dimensionality reduction, and the root dataset’s data distribution on our detection method’s performance. We repeat each experiment three times and average the results. All experiments run on a cloud server with eight NVIDIA RTX 2080 GPUs.
On Fashion-MNIST and HAR datasets, SkyMask improves the testing accuracy by 1–3% compared with the best result of all five Byzantine-robust algorithms under the same attacks. More significantly, our method gains a 10% improvement in testing accuracy on CIFAR-10 dataset under Min-Max attack.
IV-A Experiment Settings
IV-A1 IID and non-IID Datasets
Following previous work [4, 9], in different datasets for classification tasks, if there are $C$ classes, we construct $C$ groups. The probability of a training sample with label $k$ assigned to group $k$ is $q>0$ and to other group is $\frac{1-q}{C-1}$. Each client randomly chooses one group and samples a certain number of samples uniformly from this group for its local dataset. The probability $q$ controls the degree of non-IID. If $q=\frac{1}{C}$, the dataset for each client is IID. For the root dataset, we randomly sample a certain number of samples from the training dataset with a bias probability $p$ whose meaning is similar to $q$ and then split the rest for each client in the way mentioned above. We use Fashion-MNIST, CIFAR-10, and HAR dataset [22] as the evaluation datasets, which have been widely used in prior FL studies. The Fashion-MNIST dataset consists of a 60K image samples training set and a 10K samples test set, where each image is associated with a label from 10 classes. The CIFAR-10 dataset composes 50K training images and 10K test images in 10 classes. They are colorful with three channels, contain many features noise, and have varying object scales. For both Fashion-MNIST and CIFAR-10, we use $q=0.5$ to conduct the non-IID experiments as [4]. By setting $q=0.1$, we also conduct the IID experiments as [12].
HAR dataset is built by recording 30 participants’ daily living with sensors, consisting of 7,352 pieces of 561-feature vectors with time and frequency domain variables. The main task is to classify the activity into one of six human activities. Since they are collected naturally by 30 devices, it is unnecessary to divide the dataset manually. We randomly assign 70% of each device’s dataset as the training dataset and 30% to be the testing data.
IV-A2 Models
@@
We construct a CNN with two convolutional layers and two linear layers as the global model for Fashion-MNIST dataset. For CIFAR-10 dataset, to verify SkyMask’s effectiveness on large models in FL, we use a more complex model, ResNet20. We select a linear regression classifier for HAR dataset.
IV-A3 Baseline attacks
The experiments cover untargeted, targeted, and backdoor attacks, including the following baselines:
  • Fang attack [4]: Fang attack maximizes the deviation of the global model towards the opposite direction of the unattacked global model update. By using Trim and Krum as the aggregation algorithm in the attack, we obtain two different attacks: Fang-Trim and Fang-Krum.
  • Label-flipped (LF) [2]: Following the same setting as Fang [4], we flip label $k$ to $C-k-1$ for each sample on malicious clients, where $k\in\{0,\dots,C-1\}$ and $C$ denotes the total number of dataset classes.
  • AGR-agnostic [12], i.e., Aggregation algorithm agnostic attack, aims to limit the range of malicious updates and maximize the perturbation. It has two variants: Min-Max attack restricts the range based on the maximum value of the L-$2$ distance between any pair of benign model updates. Min-Sum attack is based on the minimum value of the sum of squares of the distances between any benign model update and all other benign model updates.
  • Scaling attack [15]: Scaling attack is a targeted backdoor attack. It poisons the local dataset with trigger-embedded samples and scales up the model updates before sending them to the server. We use the same pattern trigger in LIE [5] and follow the same data augmentation scheme as FLTrust [9].
IV-A4 Baseline defenses
@@
In §IV-B, to show the defense effectiveness, we compare SkyMask with five Byzantine-robust FL aggregation algorithms: FLTrust [9], Trim [10], Krum, MKrum [8] and Bulyan [11]. In §IV-C, to reveal the significance of learnable masks in malicious client detection, we compare SkyMask with two malicious client detection algorithms: Tolpegin defense [2] and FLDetector [16].
@@
TABLE I: FL testing accuracy under different attacks and attack success rates of Scaling attack. The experimental results of Scaling attack is in the form of “testing accuracy / attack success rate”.
@@
Dataset (Model) Attack FedAvg FLTrust Trim Krum MKrum Bulyan SkyMask-NR SkyMask
Fashion -MNIST (CNN) non-IID None 0.89 0.89 0.88 0.83 0.88 0.87 0.89 0.89
@@
Fashion -MNIST (CNN) IID None 0.90 0.89 0.89 0.86 0.89 0.87 0.89 0.90
LF 0.81 0.88 0.84 0.86 0.89 0.87 0.89 0.90
Min-Max 0.63 0.89 0.76 0.86 0.59 0.72 0.90 0.90
Min-Sum 0.85 0.89 0.78 0.55 0.84 0.68 0.90 0.90
Fang-Trim 0.53 0.89 0.74 0.86 0.84 0.86 0.90 0.90
Fang-Krum 0.88 0.89 0.86 0.26 0.71 0.78 0.90 0.90
Scaling 0.83 / 0.21 0.89 / 0.10 0.88 / 0.14 0.85 / 0.11 0.89 / 0.11 0.87 / 0.11 0.89 / 0.10 0.90 / 0.10
CIFAR-10 (ResNet20) non-IID None 0.77 0.75 0.77 0.54 0.72 0.72 0.76 0.76
@@
CIFAR-10 (ResNet20) IID None 0.79 0.78 0.78 0.63 0.73 0.73 0.79 0.79
LF 0.76 0.76 0.74 0.64 0.72 0.73 0.77 0.78
Min-Max 0.68 0.61 0.58 0.56 0.65 0.69 0.79 0.79
Min-Sum 0.73 0.75 0.62 0.30 0.66 0.70 0.79 0.79
Fang-Trim 0.10 0.70 0.38 0.62 0.54 0.69 0.78 0.78
Fang-Krum 0.68 0.77 0.60 0.42 0.34 0.34 0.79 0.79
Scaling 0.10 / 1.00 0.76 / 0.12 0.75 / 0.67 0.62 / 0.10 0.74 / 0.10 0.74 / 0.10 0.78 / 0.11 0.79 / 0.10
HAR (LR) None 0.96 0.96 0.96 0.81 0.95 0.94 0.96 0.96
LF 0.88 0.95 0.81 0.86 0.79 0.74 0.96 0.96
Min-Max 0.89 0.95 0.93 0.76 0.83 0.66 0.96 0.96
Min-Sum 0.96 0.95 0.94 0.30 0.86 0.83 0.96 0.96
Fang-Trim 0.23 0.96 0.80 0.77 0.65 0.82 0.96 0.96
Fang-Krum 0.96 0.96 0.93 0.15 0.75 0.71 0.96 0.96
Scaling 0.72 / 0.20 0.96 / 0.20 0.92 / 0.16 0.84 / 0.19 0.96 / 0.20 0.94 / 0.19 0.96 / 0.19 0.96 / 0.20
IV-A5 FL parameter settings
We use 100 clients in Fashion-MNIST and CIFAR-10 experiments, and in HAR, we use 30 clients, following previous study [9]. To boost the effectiveness of attacks, we apply multiple-client attacks instead of single-client attacks, and the default fraction of malicious clients is set to 20%. The default binarization threshold $\tau$ is $0.5$, and the root dataset bias probability $p=0.1$.
We set both global model learning rate $\alpha$ and local model learning rate $\beta$ to $0.5$ in all experiments. For mask learning rate $\gamma$, we set $1\times 10^{7}$ in Fashion-MNIST and HAR datasets but $1\times 10^{8}$ in CIFAR-10 dataset. This is because the order of magnitude of the gradient is $10^{-8}-10^{-7}$, which declines along the model’s dimensionality increases, so we should set it larger to help masks converge. In CIFAR-10, we allow the system to execute server-client communication for $500$ rounds and let clients do $l=5$ local training iterations per round. While in Fashion-MNIST and HAR, the system runs $2,500$ and $1,000$ communication rounds, respectively, and each client only does a single local training iteration per round.
In our experiments, the root dataset has 100 samples and no intersection with the training and test datasets. This setting is similar to FLTrust [9]. 100 is enough for our defense, and it is easy for the server to collect. So we do not discuss the influence of the size of the root dataset.
IV-A6 SkyMask variants
Intuitively, if the fraction of malicious clients is less than 50% in the real world, we can choose the larger one in two clusters of masks as benign. This way, we can further reduce the computational pressure on the server. To verify the effectiveness with or without the trusted root model, we consider the following two schemes: SkyMask-NR (no trusted root model): The FL system does not generate a trusted root model, and after all the masks are converged, the system chooses the larger one in two clusters of the masks as benign for aggregation. SkyMask: The FL system uses the defense strategy fully following the default steps described in §III-C.
IV-A7 Metrics
Our experiments compare the testing accuracy of global models of various FL defenses under different attacks, obtained from the prediction accuracy of the global model on the test dataset. The attack success rate under Scaling attack is the fraction of the samples in the test dataset that are identified as the target label after being embedded with triggers.
To show the malicious client detection ability, we also record false positive rate (FPR) and false negative rate (FNR) in the malicious client detection stage. FPR indicates the number of benign clients that are misidentified as malicious clients divided by the total number of benign clients, and FNR indicates the number of malicious clients that are misidentified as benign clients divided by the total number of malicious clients. The Mean FPR/FNR is the average of FPR/FNRs in each communication round.
IV-B The Defense Effectiveness of SkyMask
We first certificate the normal defensive capability under a low fraction of attacks. Then, there is also possibly no attacker in the FL system, so the defense methods should not affect the global model’s performance. Eventually, we consider an extreme condition that the fraction of malicious clients is high.
Figure 4: The impact of high fractions of malicious clients.
IV-B1 Under a low fraction of attacks
@@
By default, we set the fraction of malicious clients to 20% following the default setting in [9]. Table I shows that, on Fashion-MNIST dataset, Multi-Krum and Bulyan algorithms perform robustness under LF attack, but they fail in the other attacks. FLTrust is relatively good that in a few experiments it achieves the same performance as ours in the non-IID case, while under LF attack, it has a 3% accuracy loss. Our SkyMask achieves the highest accuracy rate on Fashion-MNIST under different attacks compared to all others. Scaling attack does not show an effective impact on all the defense methods.
@@
To demonstrate the generalization, we conduct the experiments on CIFAR-10 dataset using a larger model, ResNet20. In the IID and non-IID cases, FLTrust does not perform as well as it does on FashionMNIST dataset, with a severe drop in accuracy under multiple attacks (especially the Min-Max attack). The other defense methods have little or no defensive effect. On the contrary, our SkyMask reaches the highest accuracy under attacks. In the IID case, our method achieves the most significant gap with other algorithms in Min-Max attack experiments. The testing accuracy is up to $0.79$, while the accuracy results obtained by others are less than $0.69$.
Since the samples with target label 0 are 10% of the test dataset in FashionMNIST and CIFAR-10, the attack success rate is about 10% if there is no attack. We can see that FedAvg is overwhelmed and Trim suffers from the Scaling attack severely on CIFAR-10. SkyMask still keeps the robustness and performs well.
The HAR dataset has a real-world data distribution, making the experiment more meaningful and closer to reality. We can see in the HAR part of Table I that Trim, Krum, and their variants cannot work well. FLTrust always has a little gap with the accuracy of models not under attack, which means it could not eliminate the impact of attacks, but it is still acceptable. Unlike all these defense methods, our SkyMask keeps the global model immune to attacks.
@@
Due to the attack designers’ optimization for the abnormality degree, only some particular parameters of the poisoning model updates changed with emphasis. It can be seen that the four fine-grained model poisoning attacks (Min-max, Min-Sum, Fang-Trim, and Fang-Krum) bypass the existing defense methods in many cases.
@@
As Section IV-A6 describes, the server does not need to consume resources to generate a root model if the fraction of malicious clients is less than 50%. In the 20% case, as shown in Table I, SkyMask-NR achieves good results under any attack and on any dataset, which shows the same defensive capability as SkyMask. Therefore, if the server confirms that the number of malicious clients is less than half the total number, it can choose SkyMask-NR as the defense strategy to reduce the overhead while maintaining the same defensive capability. For generalization, SkyMask can be applied to handle more cases.
IV-B2 Under no attack
@@
Some attacks send poisoned models to the server only at certain communication rounds, and some model updates are often mistaken for poisoning by existing defense methods due to data heterogeneity. The system collapses if the global model performance suffers from the defense strategy, even when there is no attack. So when there is no attack, the expected result is that any method should not affect the performance of the basic aggregation algorithm FedAvg can achieve.
In Table I, almost all other robust aggregation algorithms affect the convergence prediction accuracy in the no-attack case. For instance, FLTrust has an accuracy loss of about 2% for the non-IID case on CIFAR-10 dataset, and Krum algorithm has an accuracy loss ranging from about 4% to 20% on different datasets.
@@
At the same time, our SkyMask maintains a comparable performance as unattacked FedAvg on all datasets with a difference of less than 1%. The existing robust aggregation algorithms either choose a subset of model updates that seem benign or use model updates’ statistics to correct the impact. Hence, it is possible to introduce harmful impacts without attack. If there is no attack, our masks form only one cluster, or there are only several outliers. So the server selects most clients and has the same or very similar convergence with FedAvg in both IID and non-IID cases.
@@
TABLE II: Testing accuracy, FPR and FNR of different malicious client detection methods under different attacks. The experimental results of Scaling attack is in the form of “testing accuracy / attack success rate”.
@@
Fashion -MNIST (CNN) non-IID None 0.89 0.86 0.89 / / / / / /
LF 0.89 0.86 0.89 0.11% 0.05% 0.00% 0.38% 0.00% 0.02%
Min-Max 0.88 0.08 0.89 0.38% 96.94% 0.27% 0.80% 100% 0.40%
Min-Sum 0.64 0.38 0.89 41.1% 100% 0.26% 84.8% 100% 0.00%
Fang-Trim 0.88 0.86 0.89 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
Fang-Krum 0.85 0.10 0.89 28.2% 99.46% 0.20% 60.4% 100% 0.00%
Scaling 0.89 / 0.10 0.89 / 0.10 0.89 / 0.10 0.00% 0.00% 0.00% 0.03% 0.00% 0.00%
Fashion -MNIST (CNN) IID None 0.90 0.88 0.90 / / / / / /
LF 0.89 0.87 0.90 0.07% 0.02% 0.00% 0.00% 0.00% 0.00%
Min-Max 0.89 0.87 0.90 0.06% 0.00% 0.00% 0.00% 0.81% 0.00%
Min-Sum 0.88 0.49 0.90 15.34% 100% 0.01% 7.60% 100% 0.00%
Fang-Trim 0.90 0.88 0.90 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
Fang-Krum 0.89 0.13 0.90 8.04% 99.60% 0.19% 6.40% 100% 0.00%
Scaling 0.89 / 0.11 0.90 / 0.11 0.90 / 0.10 0.00% 0.00% 0.00% 0.68% 0.00% 0.00%
CIFAR10 (ResNet20) non-IID None 0.76 0.75 0.76 / / / / / /

### v2

@@ 3.3 Why SkyMask Works
Figure 5: The visualization of the training process of masks applied to malicious and benign models.
@@
We choose a four-layer CNN model and Fashion-MNIST dataset, and set 20 of 100 clients as malicious. Under Fang-Trim attack, we take the mask training process in one communication round. We calculate the proportion of “1”s in each part of the binary mask. A smaller value means a fewer parameters of this client are selected for the aggregated masked model. For simplicity, here only shows 20 masks for malicious and 20 for benign clients in each mask training iteration.
@@
As shown in Fig. 5, during the mask training process, the binary parameters in the “conv1.weight”, “conv2.weight”, and “fc1.weight” parts of malicious model updates are set to zeros, which are more often than those of benign updates. It is related to the fact that the attack prefers to poison the parameters in specific parts.
The masks gradually reduce the poisoning effect of the toxic parameters during the training so that we can classify malicious clients with the help of the masks. Therefore, SkyMask can perceive the poisoning attacks at the parameter level in the model updates without knowing the details of the attack, which makes it attack-agnostic.
3.4 SkyMask’s Complexity
@@
Let $O(T)$ denote the computational complexity of one local training iteration, and the model parameters’ dimensionality is $V$.
@@
The computational complexities of existing defense methods are as follows: $O(nV)$ for FLTrust, $O(nV\log n)$ for Trim, $O(n^{2}V)$ for Krum, $O(T+nV)$ for DeepSight, and $O(nV)$ for FLAME. For SkyMask, to update the masks once, the server computes the gradient of each mask from the gradient of the aggregated masked model $\tilde{W}_{t+1}$ by backpropagation, with a computational complexity of $O(nV)$. Hence, the total computational complexity of SkyMask is $O((T+nV)t_{m})+O(nV)=O(Tt_{m}+nVt_{m})$, where $t_{m}$ is the number of mask training convergence iterations. Since there is no data dependency between any mask parameters, we can use parallel training [6] to accelerate the mask optimization process so that the time overhead can be further reduced, trading computation power for time. The above analysis shows that the additional cost of the mask learning mechanism over other defense methods is equivalent to the cost of conducting several iterations of local training, which is acceptable compared to the computational power of the server.
@@
The space complexity of existing defense methods are: $O(nV)$ for Trim, $O(n+nV)$ for FLTrust, $O(n^{2}+nV)$ for Krum, DeepSight, and FLAME. For SkyMask, the additional space complexity is proportional to the number of clients $n$ and the space complexity of the global model parameters because SkyMask assigns a mask for each client and the masks have the same dimensionality as the global model. Considering that the server stores $n$ local models’ parameters in the current round, the total space cost is about twice the others, and its space complexity is $O(nV)$, which is reasonable.
4 Experiments
@@
We evaluate our method with existing Byzantine-robust FL aggregation algorithms and two malicious client detection algorithms. To verify the generalization of our method, we conduct experiments using various popular datasets paired with different models. We evaluate our method with seven popular attacks, explore the impact of the fraction of malicious clients, and verify its scalability. In sensitivity analysis, we discuss the influence of binarization threshold, the number of features after dimensionality reduction, and root dataset’s data distribution on our method’s performance (details in supplementary §C), respectively.
4.1 Experiment Settings
4.1.1 Datasets and models.
@@ 4.1 Experiment Settings
We evaluate SkyMask on three widely-used datasets: Fashion-MNIST, CIFAR-10, and CIFAR-100 [13]. Following previous methods [7, 5], we construct non-IID local dataset with bias probability $p=0.5$ for each client and the root dataset for the server. We construct a CNN with two convolutional layers and two linear layers as the global model for Fashion-MNIST dataset. For CIFAR-10 and CIFAR-100 datasets, to verify SkyMask’s effectiveness on large models in FL, we use a more complex model, ResNet20.
4.1.2 Baseline attacks and defenses.
@@
The experiments cover 1) untargeted attacks: Label-flipped (LF) [30] and four fine-grained attacks, including Fang attacks [7] (Fang-Trim and Fang-Krum attacks) and two types of the AGR-agnostic attack [28] (Min-Max and Min-Sum attacks); 2) targeted attack: Scaling attack [1] and DBA [32]. In §4.2, to show the defense effectiveness, we compare SkyMask with five Byzantine-robust FL aggregation algorithms: FLTrust [5], Trim [33], Krum [4], DeepSight [27] and FLAME [24]. In §4.3, to reveal the significance of learnable masks in malicious client detection, we compare SkyMask with two malicious client detection algorithms: Tolpegin defense [30] and FLDetector [36].
4.1.3 FL parameter settings.
Following the previous study [5], we use 100 clients in all experiments and apply multiple-client attacks. The default fraction of malicious clients is set to 20%. In our experiments, the root dataset has 100 samples and no intersection with the training and test datasets, which is similar to FLTrust [5]. Other hyperparameters are detailed in supplementary §A.4.
4.1.4 SkyMask variants.
Intuitively, if the fraction of malicious clients is less than 50% in the real world, we can choose the larger one in two clusters of masks as benign. This way, we can further reduce the computational pressure on the server. To verify the effectiveness with or without the trusted root model, we consider the following two schemes: 1) SkyMask-NR (no trusted root model). The FL system does not generate a trusted root model, and after all the masks are converged, the system chooses the larger one in two clusters of the masks as benign for aggregation; 2) SkyMask. The FL system uses the defense strategy fully following the default steps described in §3.2.
4.1.5 Metrics.
The testing accuracy is obtained from the prediction accuracy of the global model on the test dataset. The attack success rate under targeted attack is the fraction of the trigger-embedded samples in the test dataset that are identified as the target label. To show the malicious client detection ability, we record false positive rate (FPR) and false negative rate (FNR) in the malicious client detection stage. FPR indicates the number of benign clients that are misidentified as malicious clients divided by the total number of benign clients, and FNR indicates the number of malicious clients that are misidentified as benign clients divided by the total number of malicious clients. The Mean FPR/FNR is the average of FPR/FNRs in each communication round.
4.2 The Defense Effectiveness of SkyMask
We first certificate the defensive capability under a low fraction of attacks. Then, there is also possibly no attacker in the FL system, so the defense methods should not affect the global model’s performance. Eventually, we consider an extreme condition that the fraction of malicious clients is high.
@@ 4.2 The Defense Effectiveness of SkyMask
Table 1: FL testing accuracy under different attacks and attack success rates of targeted attack. The experimental results of targeted attacks are in the form of “testing accuracy/attack success rate.”
@@
Dataset (Model) Attack FedAvg FLTrust Trim Krum DeepSight FLAME SkyMask-NR SkyMask
Fashion -MNIST (CNN) None 0.89 0.89 0.88 0.83 0.88 0.87 0.89 0.89
@@
DBA 0.86/0.51 0.88/0.18 0.88/0.31 0.83/0.11 0.89/0.16 0.89/0.11 0.89/0.11 0.89/0.10
CIFAR-10 (ResNet20) None 0.77 0.75 0.77 0.54 0.72 0.72 0.76 0.76
@@
DBA 0.72/0.99 0.76/0.88 0.76/0.97 0.56/0.09 0.77/0.16 0.77/0.14 0.77/0.11 0.77/0.10
CIFAR-100 (ResNet20) None 0.44 0.39 0.43 0.17 0.44 0.44 0.44 0.44
LF 0.41 0.39 0.38 0.11 0.43 0.43 0.44 0.44
Min-Max 0.16 0.30 0.16 0.05 0.16 0.19 0.44 0.44
Min-Sum 0.33 0.28 0.33 0.17 0.35 0.34 0.44 0.44
Fang-Trim 0.01 0.34 0.04 0.15 0.20 0.22 0.44 0.44
Fang-Krum 0.03 0.37 0.04 0.03 0.03 0.02 0.44 0.44
Scaling 0.01/1.00 0.36/0.30 0.43/0.98 0.14/0.00 0.44/0.04 0.44/0.08 0.44/0.01 0.44/0.01
DBA 0.37/0.98 0.40/0.90 0.44/0.94 0.16/0.01 0.45/0.16 0.45/0.15 0.44/0.02 0.44/0.01
4.2.1 Under a low fraction of attacks.
@@
By default, we set the fraction of malicious clients to 20% following the default setting in [5]. Table 1 shows that the testing accuracy of most existing defenses is reduced by more than 10% with at least one attack. FLTrust is a relatively strong baseline defense because it achieves similar good performance as ours in a few Fashion-MNIST experiments. While under LF attack, FLTrust has a 3% accuracy loss. Our SkyMask achieves the highest testing accuracy on Fashion-MNIST under different attacks compared to all others.
@@
To demonstrate the generalization, we experiment on CIFAR-10 and CIFAR-100 dataset using a larger model, ResNet20. In such scenarios, multi-client fine-grained attacks are much more powerful. Trim, Krum, DeepSight, and FLAME are out of defense. FLTrust does not perform as well as it does on Fashion-MNIST dataset, with a severe drop in accuracy under various attacks. In contrast, our SkyMask reaches the highest accuracy under all attacks. Especially on CIFAR-100 dataset, our method achieves the most significant gap with other algorithms in Min-Max attack experiments. The accuracy is up to $0.44$, while the accuracy results obtained by others are less than $0.3$. Besides, SkyMask achieves the best and unattacked-level performance under the targeted attacks.
@@
Due to the attackers’ optimization for the abnormality degree, only some particular parameters of the poisoning models changed with emphasis. The results show that the four fine-grained attacks (Min-max, Min-Sum, Fang-Trim, and Fang-Krum attacks) bypass the existing defense methods in many cases.
@@
As §4.1.4 describes, the server does not need to consume resources to generate a root model if the fraction of malicious clients is less than 50%. In the 20% case, as shown in Table 1, SkyMask-NR achieves good results under any attack and on any dataset, which shows the same defensive capability as SkyMask. Therefore, if the server confirms that the number of malicious clients is less than half of the total number, it can choose SkyMask-NR as the defense strategy to reduce the overhead while maintaining the same defensive capability. For a more general scenario, it can choose SkyMask as the defense strategy.
4.2.2 Under no attack.
@@
Some attacks send poisoned models to the server only at certain rounds, and some model updates can be mistaken for malicious by existing defense methods due to data heterogeneity. The system collapses if the global model performance suffers from the defense, even when there is no attack. So when there is no attack, the expected result is that any method should not affect the performance achieved by the basic aggregation algorithm FedAvg.
@@
In Table 1, almost all the existing robust aggregation algorithms affect the testing accuracy in the no-attack case. For instance, FLTrust has an accuracy loss of about 2% on CIFAR-10 dataset and 5% on CIFAR-100. Krum algorithm has an accuracy loss ranging from about 6% to 27% on different datasets. In contrast, our SkyMask maintains a comparable performance as unattacked FedAvg on all datasets with a difference of less than 1%. The existing robust aggregation algorithms either choose a subset of model updates that seem benign or use model updates’ statistics to correct the impact. Hence, it is possible to introduce harmful impacts without attack. If there is no attack, our masks form only one cluster, or there are only several outliers. So SkyMask selects most of clients and has the same or very similar convergence compared with FedAvg.
Figure 6: The impact of high fractions of malicious clients under fine-grained attacks.
4.2.3 Under a high fraction of attacks.
The experiments mentioned in the previous sections set the fraction of malicious clients to 20%, while in a real application environment, an attacker can control a larger fraction of malicious clients for the attack. Therefore, to verify the defensive capability of our system in this case, we use four-layer CNN as the global model on Fashion-MNIST dataset and conduct comparison experiments with 40%, 60%, and 80% malicious fractions. We present the results pertaining to fine-grained attacks in Fig. 6. Additional results can be found in the supplementary §B.1.
Only SkyMask and FLTrust maintain the defensive capability under fine-grained attacks. The other defense algorithms completely lose their defensive capability, exhibiting a significant gap from the unattacked level. From Fig. 6(b), Fig. 6(c), and Fig. 6(d), we can see that SkyMask converges more stable than FLTrust when they are attacked by Fang-Krum, Min-Max, and Min-Sum attacks. In Fig. 6(d), the accuracy obtained by SkyMask is 2% higher than FLTrust’s under Min-Sum attacks.
When the fraction is more than 50%, only SkyMask maintains the defensive capability. Even as the best-performing baseline defense, FLTrust’s performance worsens as the fraction rises. FLTrust’s testing accuracy fluctuates more, and the global model even converges in the wrong direction, e.g. in Fig. 6(l), the accuracy obtained by FLTrust decreases from 85% to 81% after a sharp fluctuation. Under other attacks, the increasing fraction influences FLTrust’s convergence speed, whereas SkyMask remains unaffected and converges faster than FLTrust.
@@
Table 2: Testing accuracy, FPR, and FNR of different malicious client detection methods under different attacks. The experimental results of targeted attacks are in the form of “testing accuracy/attack success rate.”
@@
CIFAR10 (ResNet20) None 0.76 0.75 0.76 0.00% 18.2% 0.00% / / /